# pFedMMA: Personalized Federated Fine-Tuning with Multi-Modal Adapter for Vision-Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) like CLIP have demonstrated remarkable generalization in zero- and few-shot settings, but adapting them efficiently to decentralized, heterogeneous data remains a challenge. While prompt tuning has emerged as a popular parameter-efficient approach in personalized federated learning, existing methods often sacrifice generalization in favor of personalization, struggling particularly on unseen classes or domains. In this work, we propose pFedMMA, the first personalized federated learning framework that leverages multi-modal adapters for vision-language tasks. Each adapter contains modality-specific up- and down-projection layers alongside a globally shared projection that aligns cross-modal features. Our asymmetric optimization strategy allows clients to locally adapt to personalized data distributions while collaboratively training the shared projection to improve global generalization. This design is also communication-efficient, as only the shared component is exchanged during communication rounds. Through extensive experiments across eleven datasets, including domain- and label-shift scenarios, we show that pFedMMA achieves state-of-the-art trade-offs between personalization and generalization, outperforming recent federated prompt tuning methods. The code is available at https://github.com/sajjad-ucsb/pFedMMA.
@@ 1 Introduction
Vision-Language Models (VLMs) like CLIP [1] have revolutionized multi-modal learning by jointly embedding visual and textual data through massive contrastive pre-training [2, 3, 4]. This paradigm empowers models to generalize effectively in zero-shot and few-shot settings [5, 6, 7, 8]. Among them, larger transformer-based variants [9] (e.g., CLIP ViT-L/14) consistently outperform smaller counterparts such as ViT-B/16, with margins exceeding 6% on benchmarks like ImageNet [10]. However, the computational demands of fine-tuning such large-scale models with billions of parameters pose significant challenges, particularly for domain-specific tasks [11]. To mitigate this, Parameter-Efficient Fine-Tuning (PEFT) techniques have emerged, especially in NLP. These methods, including adapters [12, 13, 14] and prompt tuning [15, 16], introduce a lightweight set of trainable parameters or tokens, allowing the backbone model to remain frozen.
@@
While highly effective in centralized settings, these techniques fall short in scenarios involving decentralized and privacy-sensitive data, such as healthcare, legal, or industrial domains [17, 18, 19]. Federated Learning (FL) offers a promising alternative by enabling collaborative training without raw data sharing. In FL, clients independently update their local models and transmit only intermediate model updates such as parameters or gradients, which are aggregated into a global model by a central server [20].
@@
In real-world scenarios, client data often exhibits variations in domain discrepancies (feature shift) [21] or imbalanced class distributions (label shift) [22]. Simply applying standard aggregation strategies such as FedAvg [20] over prompts [23] or other fine-tuning methods such as LoRA often leads to suboptimal performance due to data heterogeneity [24, 25, 26]. As a result, Personalized Federated Learning (PFL), particularly with prompt tuning, has gained increasing attention. pFedPrompt [27] introduces personalization by coupling a global text prompt with local visual attention modules to tailor predictions to each client’s data, FedOTP [28] uses Optimal Transport to align local and global representations under label shift, and FedPGP [29] applies prompt-wise contrastive learning to enhance inter-client generalization. Recently, pFedMoAP [30] proposes a Mixture-of-Experts framework, where prompts from other clients serve as non-local experts, and each client learns an attention-based gating mechanism for selective adaptation. While these methods achieve impressive personalization performance, they often struggle to generalize to unseen classes or domains, limiting their applicability in out-of-distribution scenarios.
@@
Beyond prompt tuning, adapters offer another PEFT strategy by introducing small trainable modules into frozen pre-trained models [31, 12, 32, 33, 34, 5]. Unlike prompts, adapters operate independently of model architecture and can be easily inserted into various backbones, such as ResNets [35], ViTs [36], and Swin Transformers [37]. However, most adapter methods like AdaptFormer [12] and LoRA [34] are uni-modal and do not account for the cross-modal dependencies inherent in VLMs like CLIP [1]. Multi-modal adapters [38] address this by integrating both visual and textual signals via a shared projection layer that promotes feature alignment across modalities while preserving modality-specific knowledge. Despite their demonstrated advantages over prompt-based approaches [38, 39], their integration with PFL remains largely unexplored.
@@
In this work, we introduce a Personalized Federated Multi-Modal Adapter (pFedMMA) architecture that adopt a multi-modal adapter design with three components: a modality-specific down-projection, a shared projection, and a modality-specific up-projection. During training, all components are updated locally by each client, but only the shared projection is globally aggregated. This asymmetric training scheme enables effective personalization through client-specific projections, while promoting generalization via a shared modality-alignment space. Moreover, since only the shared adapter is communicated during rounds, the method remains communication-efficient. As validated in our experiments, this design achieves a strong trade-off between personalization and generalization under both feature and label shifts.
We summarize our main contributions as follows:
  • We observe that while most state-of-the-art prompt tuning methods achieve strong personalization performance, they often generalize poorly to unseen classes. To address this, we introduce a multi-modal adapter framework that explicitly aims to balance personalization and generalization in federated vision-language learning.
  • We propose pFedMMA, the first adapter-based approach for PFL of VLMs. Our architecture incorporates modality-specific up- and down-projection layers and a shared cross-modal projection. All components are updated locally, but only the shared projection is aggregated globally, enabling effective asymmetric optimization.
  • We conduct extensive experiments on widely used benchmarks to evaluate pFedMMA’s performance on base-to-novel generalization across both category- and domain-level tasks under heterogeneous data distributions. Results demonstrate the superiority of our approach in harmonizing generalization and personalization.
2 Related Work
2.1 Personalized Federated Learning
Personalized Federated Learning (PFL) has emerged as a pivotal research direction to address the limitations of conventional federated learning [20] when faced with heterogeneous client data. Unlike standard FL, which learns a single global model, PFL aims to produce tailored models for individual clients, thus better coping with statistical and systemic heterogeneity [40, 41]. Several personalization strategies have been proposed, including local fine-tuning [42, 40, 43], regularization-based optimization [44, 45, 46], and parameter decomposition into shared and client-specific components [47, 48, 49]. Other methods pursue clustering of clients to exploit latent similarities [50, 51, 52, 53], or leverage attention mechanisms and adaptive layers [54, 55, 56]. To further improve adaptability, techniques like FedBN [21] and PartialFed [57] address feature shift via local normalization or selective personalization. Hybrid global-local learning approaches have also been developed [58, 59]. FedOT [60] proposes learning optimal transport maps that align local distributions to a shared probability space, enabling a global classifier to be trained more effectively; personalization is achieved by composing this shared model with each client’s transport map. While these approaches have demonstrated success, they typically center on traditional ML architectures and do not yet fully leverage the potential of large pre-trained models, such as vision-language or foundation models, for personalization.
2.2 Federated Prompt Learning for VLMs
Federated Prompt Learning (FPL) extends the flexibility of prompt tuning for adapting large pre-trained models such as CLIP [1] to the FL settings, enabling efficient and personalized downstream task adaptation across decentralized clients. Early works like CoOp [61] and CoCoOp [62] laid the foundation by introducing learnable continuous prompt vectors, which sparked interest in federated extensions. PromptFL [23] and FedPrompt [63] introduced FL-style prompt aggregation, performing FedAvg [20] over client-specific prompt updates. FedPR [64] explores visual prompt learning within the null space of global prompts for MRI reconstruction, while FedAPT [65] focuses on domain-adaptive prompt tuning for cross-domain image classification. To enhance personalization, pFedPrompt [27] introduces a non-parametric attention module over local few-shot memory, and pFedPG [66] and FedTPG [67] design server-side prompt generators to issue personalized prompts to each client. FedCLIP [68] integrates attention-based adapters to better exploit the pre-trained model’s knowledge. Furthermore, FedOTP [28] leverages Optimal Transport to align global and local prompts, and FedPGP [69] utilizes prompt-wise contrastive losses to better capture diverse category-level traits across clients. Recently, pFedMoAP [30] rethinks prompt sharing by treating pre-aggregated prompts from other clients as non-local experts in a Mixture-of-Experts framework, enabling effective personalization via a lightweight, attention-based gating mechanism. Theoretical analysis of FPL [70] provides deeper understanding of its convergence properties.
2.3 Efficient Transfer Learning for VLMs
Traditional transfer learning approaches typically fine-tune all parameters of pre-trained VLMs [71, 35], but this becomes increasingly impractical as model sizes scale up, especially under computational or data constraints. To mitigate this, the community has embraced parameter-efficient transfer learning strategies that modify only a small fraction of model parameters. Among these, prompt learning techniques, briefly introduced in the previous section, optimize lightweight vectors or tokens to steer the model without altering its backbone  [61, 62, 72, 73]. Although effective, they are often limited in their expressiveness or modality interaction. As a result, adapter-based methods have emerged as a powerful alternative. CLIP-Adapter [33] and Tip-Adapter [5] inject lightweight MLP layers after the image encoder to refine visual representations. Tip-Adapter further improves efficiency by caching training features for fast inference. However, these image-only approaches neglect the cross-modal nature of VLMs. To address this, MMA  [38] introduces a multi-modal adapter architecture that fuses features across the vision and language branches via a shared representation space, enabling gradient flow between modalities. Similarly, other works explore deeper adapter integration, such as inserting adapters within self-attention and MLP blocks  [74], allowing more granular control over the representation learning process. These advances mark a shift from single-stream to multi-stream adaptation, aligning with the unique demands of multi-modal tasks. In federated settings, where full model updates are prohibitive, adapter-based techniques offer a compelling balance between personalization, generalization, and communication efficiency—making them well-suited foundations for multi-model federated frameworks like ours.
@@ 3.1 Personalized Federated Learning
Traditional federated learning frameworks are designed around the principle of global consensus, where the goal is to collaboratively train a single model that generalizes well across a federation of clients. The canonical approach, FedAvg [20], formalizes this as the minimization of a weighted average of local objectives:
$\min_{\boldsymbol{\theta}}F(\boldsymbol{\theta})=\sum_{i=1}^{N}p_{i}F_{i}(\boldsymbol{\theta}),$ (1)
where $\boldsymbol{\theta}$ denotes the global model, $F_{i}(\cdot)$ represents the local empirical loss of client $i$, and $p_{i}=\frac{n_{i}}{n}$ scales the contribution of each client by its dataset size $n_{i}$, with $n=\sum_{i}n_{i}$. In this setup, each client’s local loss is computed as the average over its data: $\sum_{k=1}^{n_{i}}\mathcal{L}_{i}(\boldsymbol{\theta}\mid(\boldsymbol{x}_{k},y_{k}))$, where $\mathcal{L}_{i}$ is the local loss function and $(\boldsymbol{x}_{k},y_{k})$ is the $k$-th data point on client $i$.
In contrast, personalized federated learning (PFL) challenges the one-size-fits-all paradigm by allowing each client to maintain its own model $\boldsymbol{\theta}_{i}$. This formulation acknowledges data heterogeneity and aims to tailor learning to each client’s unique distribution. The optimization objective for PFL becomes:
$\min_{\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{N}}F(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{N})=\sum_{i=1}^{N}p_{i}F_{i}(\boldsymbol{\theta}_{i}),$ (2)
@@ 3.2 Vision-Language Classification with Few-Shot Adaptation
In vision-language classification, predictions emerge from the powerful alignment between visual and textual modalities established during pretraining. Given a label set with $K$ classes, the model begins by crafting natural language prompts [75]—semantic descriptions like “a photo of a [class name]”—for each class $c_{k}$. These textual cues are passed through a frozen text encoder $\theta_{t}$, producing normalized text embeddings $\mathbf{z}_{k}^{(T)}=\theta_{t}(c_{k})\in\mathbb{R}^{d}$. In parallel, each input image $\mathbf{x}_{i}$ is processed by a visual encoder $\theta_{v}$, generating a corresponding normalized image embedding $\mathbf{z}_{i}^{(I)}=\theta_{v}(\mathbf{x}_{i})\in\mathbb{R}^{d}$. Classification then hinges on comparing the cosine similarity between these multimodal representations. The result is a set of logits transformed into class probabilities via a temperature-scaled softmax:
$p_{i,k}=\frac{\exp\left(\cos(\mathbf{z}_{i}^{(I)},\mathbf{z}_{k}^{(T)})/\gamma\right)}{\sum_{j=1}^{K}\exp\left(\cos(\mathbf{z}_{i}^{(I)},\mathbf{z}_{j}^{(T)})/\gamma\right)},$ (3)
where $\gamma$ is the temperature parameter controlling distribution sharpness. The predicted label for image $\mathbf{x}_{i}$ corresponds to the class with the highest posterior probability:
$\hat{k}=\arg\max_{k}\,p_{i,k}.$
This zero-shot classification pipeline mirrors the contrastive training strategy employed in foundational vision-language models like CLIP [1], enabling impressive generalization to novel tasks without requiring any target-domain fine-tuning.
To further tailor the model to downstream tasks, the few-shot setting introduces a small set of labeled examples per class, typically fewer than 16. With $M$ support samples per class, and ground-truth labels encoded as one-hot vectors $y_{ik}$ (where $y_{ik}=1$ if $\mathbf{x}_{i}$ belongs to class $k$, and 0 otherwise), classification proceeds identically to the zero-shot case. However, the model is now adapted by minimizing the cross-entropy loss over the labeled support set:
$\mathcal{L}_{\text{CE}}=-\frac{1}{M}\sum_{i=1}^{M}\sum_{k=1}^{K}y_{ik}\ln p_{i,k}.$ (4)
This fine-tuning step enables the model to better capture domain-specific semantics while maintaining the efficiency and generalization capabilities of the pretrained architecture. Adaptation can be achieved through various strategies. One approach is to directly optimize the input prompts $\{c_{k}\}_{k=1}^{K}$, following the principles of prompt tuning [76]. Alternatively, lightweight task-specific modules such as adapter layers [33] or low-rank parameterizations like LoRA [77] can be fine-tuned, while keeping the backbone encoders frozen.
@@ 3.3 Fine-Tuning via Parallel Adapters
In contrast to the serial adapter architecture introduced by [78], where adapter modules are inserted sequentially after each sub-layer (e.g., attention or feed-forward), parallel adapters [79] adopt an alternative integration strategy. Rather than placing the adapter transformation after the main layer, the parallel formulation processes the input through the adapter module concurrently with the frozen backbone transformation and combines their outputs additively.
Let $x\in\mathbb{R}^{d}$ be the input to a transformer sub-layer, and let $f(x)$ denote the frozen pre-trained transformation. A parallel adapter layer computes the output as:
$\text{Output}(x)=f(x)+\alpha A(x),$ (5)
where $\alpha$ is a scaling factor and the adapter module $A(x)$ uses the same bottleneck structure as in the serial configuration:
$A(x)=U(\delta(D(x))),$ (6)
where $U$ is an up-projection affine map, $D$ is a down-projection affine map, and $\delta$ is a non-linear activation function such as ReLU. If the input $x$ has dimensionality $d$, then $D\in\mathbb{R}^{r\times d}$ and $U\in\mathbb{R}^{d\times r}$, where $r\ll d$. This bottleneck structure introduces significantly fewer trainable parameters compared to the full model. As with serial adapters, only the adapter parameters are trained during fine-tuning; the base model remains frozen. Parallel adapters preserve model expressiveness while enabling efficient adaptation with minimal architectural modifications.

### v2

@@ Abstract
Vision-Language Models (VLMs) like CLIP have demonstrated remarkable generalization in zero- and few-shot settings, but adapting them efficiently to decentralized, heterogeneous data remains a challenge. While prompt tuning has emerged as a popular parameter-efficient approach in personalized federated learning, existing methods often sacrifice generalization in favor of personalization, struggling particularly on unseen classes or domains. In this work, we propose pFedMMA, the first personalized federated learning framework that leverages multi-modal adapters for vision-language tasks. Each adapter contains modality-specific up- and down-projection layers alongside a globally shared projection that aligns cross-modal features. Our optimization strategy allows clients to locally adapt to personalized data distributions while collaboratively training the shared projection to improve global generalization. This design is also communication-efficient, as only the shared component is exchanged during communication rounds. Through extensive experiments across eleven datasets, including domain- and label-shift scenarios, we show that pFedMMA achieves state-of-the-art trade-offs between personalization and generalization, outperforming recent federated prompt tuning methods.
@@ 1 Introduction
Vision-Language Models (VLMs) like CLIP Radford et al. (2021) have revolutionized multi-modal learning by jointly embedding visual and textual data through massive contrastive pre-training Jia et al. (2021); Li et al. (2022); Yao et al. . This paradigm empowers models to generalize effectively in zero-shot and few-shot settings Zhang et al. (2022); Zhu et al. (2023); Ghiasvand et al. (2025); Aghdam & Hu (2025). Among them, larger transformer-based variants Vaswani (2017) (e.g., CLIP ViT-L/14) consistently outperform smaller counterparts such as ViT-B/16, with margins exceeding 6% on benchmarks like ImageNet Deng et al. (2009). However, the computational demands of fine-tuning such large-scale models with billions of parameters pose significant challenges, particularly for domain-specific tasks Oskouie et al. (2025). To mitigate this, Parameter-Efficient Fine-Tuning (PEFT) techniques have emerged, especially in NLP. These methods, including adapters Chen et al. (2022); Karimi Mahabadi et al. (2021); Rebuffi et al. (2017) and prompt tuning Jia et al. (2022); Li & Liang (2021), introduce a lightweight set of trainable parameters or tokens, allowing the backbone model to remain frozen.
@@
While highly effective in centralized settings, these techniques fall short in scenarios involving decentralized and privacy-sensitive data, such as healthcare, legal, or industrial domains Manoel et al. (2023); Shoham & Rappoport (2023); Mahjourian & Nguyen (2025). Federated Learning (FL) offers a promising alternative by enabling collaborative training without raw data sharing. In FL, clients update their local models and transmit only intermediate model updates such as parameters or gradients, which are aggregated into a global model by a central server McMahan et al. (2017).
@@
In real-world scenarios, client data often exhibits variations in domain discrepancies (feature shift) Li et al. or imbalanced class distributions (label shift) Li et al. (2021a). Simply applying standard aggregation strategies, such as FedAvg McMahan et al. (2017), over prompts Guo et al. (2023b) or other fine-tuning methods, such as LoRA, often leads to suboptimal performance due to data heterogeneity Zhang et al. (2023); Borazjani et al. (2025). As a result, Personalized Federated Learning (PFL), particularly with prompt tuning, has gained increasing attention. pFedPrompt Guo et al. (2023a) introduces personalization by coupling a global text prompt with local visual attention modules to tailor predictions to each client’s data. FedOTP Li et al. (2024) uses Optimal Transport to align local and global representations under label shift. FedPGP Cui et al. applies prompt-wise contrastive learning to enhance inter-client generalization. Recently, pFedMoAP Luo et al. (2025) proposes a Mixture-of-Experts framework, where prompts from other clients serve as non-local experts, and each client learns an attention-based gating mechanism for selective adaptation. While these methods achieve impressive personalization performance, they often struggle to generalize to unseen classes or domains, limiting their applicability in out-of-distribution scenarios. For example, as shown in Fig. 1, FedOTP achieves poor harmonic mean accuracy, even though it has been shown to have strong personalization performance.
Figure 1: Few-shot performance across datasets using the ViT-B/16 model. Each radar chart illustrates accuracy (%) for a fixed shot count, with spokes representing the evaluation datasets. Curves correspond to different methods, and values increase outward (20–100%). Accuracy is reported as the harmonic mean (HM) over local, base, and novel classes for each dataset and shot.
@@
Beyond prompt tuning, adapters offer another PEFT strategy by introducing small trainable modules into frozen pre-trained models Cai et al. (2020); Chen et al. (2022; b); Gao et al. (2024); Hu et al. (2021); Zhang et al. (2022). Unlike prompts, adapters operate independently of model architecture and can be easily inserted into various backbones, such as ResNets He et al. (2016), ViTs Dosovitskiy et al. , and Swin Transformers Liu et al. (2021). However, most adapter methods like AdaptFormer Chen et al. (2022) and LoRA Hu et al. (2021) are uni-modal and do not account for the cross-modal dependencies inherent in VLMs like CLIP Radford et al. (2021). Multi-modal adapters Yang et al. (2024) address this by integrating both visual and textual signals via a shared projection layer that promotes feature alignment across modalities while preserving modality-specific knowledge. Despite their demonstrated advantages over prompt-based approaches Yang et al. (2024); Guo & Gu (2025), their integration with PFL remains largely unexplored.
@@
In this work, we introduce a Personalized Federated Multi-Modal Adapter (pFedMMA) architecture that adopt a multi-modal adapter design with three components: a modality-specific down-projection, a shared projection, and a modality-specific up-projection. During training, all components are updated locally by each client, but only the shared projection is globally aggregated. This asymmetric training scheme enables effective personalization through client-specific projections, while promoting generalization via a shared modality-alignment space. Moreover, since only the shared adapter is communicated during rounds, the method remains communication-efficient. As confirmed by our experiments, this design achieves the strongest trade-off between personalization and generalization under both feature and label shifts. As shown in Fig. 1, on average, our proposed pFedMMA delivers the best harmonic mean performance compared to state-of-the-art federated prompt tuning methods.
Before delving into details, we summarize our contributions: (1) We observe that while most state-of-the-art prompt tuning methods achieve strong personalization performance, they often generalize poorly to unseen classes. To address this, we introduce a multi-modal adapter framework that explicitly aims to balance personalization and generalization in federated vision-language learning. (2) We propose pFedMMA, the first adapter-based approach for PFL of VLMs. Our architecture incorporates modality-specific up- and down-projection layers and a shared cross-modal projection. All components are updated locally, but only the shared projection is aggregated globally, enabling effective asymmetric optimization. (3) We conduct extensive experiments on widely used benchmarks to evaluate pFedMMA’s performance on base-to-novel generalization across both category- and domain-level tasks under heterogeneous data distributions. Results demonstrate the superiority of our approach in harmonizing generalization and personalization.
@@ 2.1 Personalized Federated Learning
Traditional federated learning frameworks are designed around the principle of global consensus, where the goal is to collaboratively train a single model that generalizes well across a federation of clients. The canonical approach, FedAvg McMahan et al. (2017), formalizes this as the minimization of a weighted average of local objectives: $\min_{\bm{\theta}}F(\bm{\theta})=\sum_{i=1}^{N}p_{i}F_{i}(\bm{\theta}),$ where $\bm{\theta}$ denotes the global model, $F_{i}(\cdot)$ represents the local empirical loss of client $i$, and $p_{i}=\frac{n_{i}}{n}$ scales the contribution of each client by its dataset size $n_{i}$, with $n=\sum_{i}n_{i}$. In this setup, each client’s local loss is computed as the average over its data: $\sum_{k=1}^{n_{i}}\mathcal{L}_{i}(\bm{\theta}\mid(\bm{x}_{k},y_{k}))$, where $\mathcal{L}_{i}$ is the local loss function and $(\bm{x}_{k},y_{k})$ is the $k$-th data point on client $i$.
In contrast, personalized federated learning (PFL) challenges the one-size-fits-all paradigm by allowing each client to maintain its own model $\bm{\theta}_{i}$. This formulation acknowledges data heterogeneity and aims to tailor learning to each client’s unique distribution. The objective for PFL becomes:
$\min_{\bm{\theta}_{1},\ldots,\bm{\theta}_{N}}F(\bm{\theta}_{1},\ldots,\bm{\theta}_{N})=\sum_{i=1}^{N}p_{i}F_{i}(\bm{\theta}_{i}),$ (1)
@@ 2.2 Vision-Language Classification with Few-Shot Adaptation
In vision-language classification, predictions emerge from the powerful alignment between visual and textual modalities established during pretraining. Given a label set with $K$ classes, the model begins by crafting natural language prompts Liu et al. (2023)—semantic descriptions like “a photo of a [class name]”—for each class $c_{k}$. These textual cues are passed through a frozen text encoder $\theta_{t}$, producing normalized text embeddings $\mathbf{z}_{k}^{(T)}=\theta_{t}(c_{k})\in\mathbb{R}^{d}$. In parallel, each input image ${\bm{x}}_{i}$ is processed by a visual encoder $\theta_{v}$, generating a corresponding normalized image embedding $\mathbf{z}_{i}^{(I)}=\theta_{v}({\bm{x}}_{i})\in\mathbb{R}^{d}$. Classification then hinges on comparing the cosine similarity between these multimodal representations. The result is a set of logits transformed into class probabilities via a temperature-scaled softmax: $p_{i,k}=\exp\left(\cos(\mathbf{z}_{i}^{(I)},\mathbf{z}_{k}^{(T)})/\gamma\right)/\sum_{j=1}^{K}\exp\left(\cos(\mathbf{z}_{i}^{(I)},\mathbf{z}_{j}^{(T)})/\gamma\right),$ where $\gamma$ is the temperature parameter controlling distribution sharpness. The predicted label for image $\mathbf{x}_{i}$ corresponds to the class with the highest posterior probability: $\hat{k}=\arg\max_{k}\,p_{i,k}.$
This zero-shot classification pipeline mirrors the contrastive training strategy employed in foundational vision-language models like CLIP Radford et al. (2021), enabling impressive generalization to novel tasks without requiring any target-domain fine-tuning.
To further tailor the model to downstream tasks, the few-shot setting introduces a small set of labeled examples per class, typically fewer than 16. With $M$ support samples per class and ground-truth labels encoded as one-hot vectors $y_{ik}$ (where $y_{ik}=1$ if $\mathbf{x}_{i}$ belongs to class $k$, and 0 otherwise), classification proceeds identically to the zero-shot case. However, the model is now adapted by minimizing the cross-entropy loss over the labeled support set: $\mathcal{L}_{\text{CE}}=-\frac{1}{M}\sum_{i=1}^{M}\sum_{k=1}^{K}y_{ik}\ln p_{i,k}.$
This fine-tuning step enables the model to better capture domain-specific semantics while maintaining the efficiency and generalization capabilities of the pretrained architecture. Adaptation can be achieved through various strategies. One approach is to directly optimize the input prompts $\{c_{k}\}_{k=1}^{K}$, following the principles of prompt tuning Chen et al. (a). Alternatively, lightweight task-specific modules such as adapter layers Gao et al. (2024) or low-rank parameterizations like LoRA Zanella & Ben Ayed (2024) can be fine-tuned, while keeping the backbone encoders frozen.
@@ 2.3 Fine-Tuning via Parallel Adapters
In contrast to the serial adapter architecture introduced by Houlsby et al. (2019), where adapter modules are inserted sequentially after each sub-layer (e.g., attention or feed-forward), parallel adapters He et al. adopt an alternative integration strategy. Rather than placing the adapter transformation after the main layer, the parallel formulation processes the input through the adapter module concurrently with the frozen backbone transformation and combines their outputs additively.
Let ${\bm{x}}\in\mathbb{R}^{d}$ be the input to a transformer sub-layer, and let $f({\bm{x}})$ denote the frozen pre-trained transformation. A parallel adapter layer computes the output as: $\text{Output}({\bm{x}})=f({\bm{x}})+\alpha{\bm{A}}({\bm{x}}),$ where $\alpha$ is a scaling factor and the adapter module $A({\bm{x}})$ uses the same bottleneck structure as in the serial configuration: ${\bm{A}}({\bm{x}})={\bm{U}}(\delta({\bm{D}}({\bm{x}}))),$ where $\bm{U}$} is an up-projection affine map, $\bm{D}$} is a down-projection affine map, and $\delta$ is a non-linear activation function such as ReLU. If the input $\bm{x}$} has dimensionality $d$, then ${\bm{D}}\in\mathbb{R}^{r\times d}$ and ${\bm{U}}\in\mathbb{R}^{d\times r}$, where $r\ll d$. This bottleneck structure introduces significantly fewer trainable parameters compared to the full model. As with serial adapters, only the adapter parameters are trained during fine-tuning, and the base model remains frozen. Parallel adapters preserve model expressiveness while enabling efficient adaptation with minimal architectural modifications.